OpenStudy (anonymous):

find the derivative of y = square root of 4/x - square root of 3x.

hero (hero):

\[y = \sqrt{\frac{4}{x}} - \sqrt{3x} \] is that it?

OpenStudy (anonymous):

yup show detailed solution...

hero (hero):

Okay, sure :)

hero (hero):

You're going to be surprised

hero (hero):

1. Simplify: \[\sqrt{\frac{4}{x}} - \sqrt{3x} = \frac{\sqrt{4}}{\sqrt{x}} - \sqrt{3x} \] \[= \frac{2}{\sqrt{x}} - \sqrt{3x} \] \[= \frac{2\sqrt{x}}{x} - \frac{x*\sqrt{3x}}{x}\] Now it's simplified a little bit for you